A woman doing seated machine curls rests her upper arm

on an arm pad, parallel to the floor as shown in Fig. 2.12. She then pulls

on the bar that is connected through a cable and a pulley to a weight

of 10 kg. The weight moves up with an acceleration of 2 m/s2. Determine

the tension in the cable.
Answer #1

Draw the free body diagram. Write the balanced force equation for the force along vertical axis. F =T-W ma=T-mg =T=m(a+g) Sub